# Infinite-dimensional next-generation reservoir computing

### v2

@@ V.4 Results
Denoting the true value $\mathbf{y}$, the estimated value $\widehat{\mathbf{y}}$, and the testing set size $h$, the following pointwise error metrics were chosen: the usual mean square error, the normalized mean square error (NMSE) given by
$\textup{NMSE}(\mathbf{y},\widehat{\mathbf{y}})=\frac{\sum_{i=1}^{h}\|\mathbf{y}_{i}-\widehat{\mathbf{y}}_{i}\|_{2}^{2}}{\sum_{i=1}^{h}\|\mathbf{y}_{i}\|^{2}_{2}},$
the mean absolute error (MAE) given by
$\textup{MAE}(\mathbf{y},\widehat{\mathbf{y}})=\frac{1}{h}\sum_{i=1}^{h}\|\mathbf{y}-\widehat{\mathbf{y}_{i}}\|_{1}$
where $\|\cdot\|_{1}$ is the 1-norm, the median absolute error (MdAE) given by
$\textup{MdAE}(\mathbf{y},\widehat{\mathbf{y}})=\frac{1}{d}\sum_{u=1}^{d}\textup{median}\left(\{|y_{u,i}-\widehat{y}_{u,i}|\}_{i=1}^{h}\right)$
where $\cdot_{u,i}$ is the $u$-th dimension of the $i$-th vector $\cdot$, the mean absolute percentage error (MAPE),
@@
for some very small $\varepsilon$, and lastly the $R^{2}$-score given by
$R^{2}(\mathbf{y},\widehat{\mathbf{y}})=\frac{1}{d}\sum_{u=1}^{d}\left(1-\frac{\sum_{i=1}^{h}(y_{u,i}-\widehat{y}_{u,i})^{2}}{\sum_{i=1}^{h}(y_{u,i}-\bar{y}_{u,i})^{2}}\right),$
where $\bar{y}_{u,i}$ denotes the average of the $u$-th dimension of the vector $\mathbf{y}$ over time steps $i=1,\ldots,h$.
Whether the estimator replicated the climate of the true time series was measured by considering the difference in the true and estimated power spectral density (PSD) and the difference in distributions using the Wasserstein-1 distance. Welch’s method with the Hann window was used to compute the PSD. The number of points per segment was chosen by visual inspection for a balance between frequency resolution and error variance. When the PSD tapers off to zero after some frequency $F$, the remaining frequencies are not considered in the final difference. Finally, the PSD error (PSDE) is computed by taking
@@
The Wasserstein-1 distance was computed to compare the distributions of the true and estimated systems. For one-dimensional systems, the Wasserstein-1 distance can be computed using scipy.stats.wasserstein_distance which uses the equivalent Cramer-1 distance, that is,
@@
where $\mu,\widehat{\mu}$ are the probability distributions of $y$ and $\widehat{y}$ respectively while CDF, $\widehat{\textup{CDF}}$ denote the cumulative distributive functions. For time series in $d$-dimensions, using scipy.stats.wasserstein_distance_nd, corresponds to solving the linear programming problem
@@
where $\Gamma(\mu,\widehat{\mu})$ is the set of probability distributions whose marginals are $\mu$ and $\widehat{\mu}$ on the first and second factors respectively. $\mu$ and $\widehat{\mu}$ are the joint distributions of $\mathbf{y}$ and $\widehat{\mathbf{y}}$ respectively. It is noted that computing the Wasserstein distance for the multidimensional case is significantly more computationally expensive than in the one-dimensional case. In the case of the Lorenz dynamical system, sampling had to be performed to make using scipy.stats.wasserstein_distance_nd tractable.
@@
The performance of the estimators was measured in the following manner for the dynamical systems. First, the valid prediction time was computed. Then, the ceiling of the best-performing valid prediction time is taken. The point-to-point error metrics MSE, NMSE, MAE, MdAE, MAPE, and $R^{2}$ are measured only up to $\lceil T_{\textup{valid}}\rceil$. Beyond the characteristic predictable timescale given by the Lyapunov time, it is not meaningful to measure the step-to-step error as the trajectories diverge exponentially according to the Lyapunov exponent. To determine if the climate is well replicated over the full testing dataset, the climate metrics PSDE and $W_{1}$ are computed for the full testing dataset, with the Lorenz needing to be sampled to compute $W_{1}$. The errors are reported in Table 3.
Table 3: Error values for each estimator and dataset. A 20% deviation is allowed for $T_{\textup{valid}}$. MSE, NMSE, MAE, MdAE, MAPE, and $R^{2}$ are computed up to $\lceil T_{\textup{valid}}\rceil$. PSDE and $W_{1}$ are computed for the full dataset (Lorenz was sampled). Except for $R^{2}$, lower is better for all errors. Scale PSDE for BEKK by $10^{-4}$.
System Estimator $T_{\textup{valid}}$ MSE NMSE MAE MdAE MAPE $R^{2}$ PSDE $W_{1}$
Lorenz NG-RC 7.178 25.910 0.252 1.827 0.0897 1.598 0.621 7.606 2.114
Polynomial 9.126 13.383 0.0960 1.155 0.0377 1.230 0.812 7.169 1.683
Volterra 10.566 6.840 0.0427 0.428 0.00385 0.222 0.907 8.325 1.982
Mackey-Glass NG-RC 0.3 0.0502 0.0548 0.188 0.174 0.235 0.0202 28.815 0.155
Polynomial 7.035 0.00358 0.00390 0.0274 0.00751 0.0329 0.930 6.831 0.00147
Volterra 8.305 0.00171 0.00186 0.0162 0.00202 0.0190 0.967 5.059 0.00138
BEKK NG-RC 0.00125 0.150 0.0204 0.0166 0.644 0.125 1.565 0.332
Polynomial 0.00124 0.150 0.0204 0.0166 0.642 0.123 1.140 0.337
Volterra 0.00102 0.136 0.0170 0.0139 0.634 0.382 0.963 0.319
@@
ct-.005inVolterra Refer to caption
@@
ct-.005inPolynomial Refer to caption
@@
ct-.005inNG-RC Refer to caption
@@
Figure 1: The PSD for the Lorenz system for the Volterra kernel is in (1(a)), the polynomial kernel is in (1(b)), and NG-RC is in (1(c)). Dimension 0 corresponds to $x$, dimension 1 to $y$, and dimension 2 to $z$. PSD for Mackey-Glass system for Volterra kernel is (1(d)), polynomial kernel is (1(e)), and NG-RC is (1(f)). Dimension 0 refers to the $z$ value. PSD for BEKK for Volterra kernel is (1(g)), polynomial kernel is (1(h)), and NG-RC is (1(i)). Only the two dimensions with the most visually prominent PSD are displayed (dimensions 84 and 117).
@@
ct-.1inVolterra Refer to caption
@@
ct-.1inPolynomial Refer to caption
@@
ct-.1inNG-RC Refer to caption
@@
Figure 2: The distributions for Lorenz system for the Volterra kernel is in (1(a)), the polynomial kernel is in (1(b)), and NG-RC is in (1(c)). The distribution for the Mackey-Glass system for Volterra kernel is (1(d)), the polynomial kernel is (1(e)), and NG-RC is (1(f)). The distributions for BEKK for Volterra kernel is (1(g)), the polynomial kernel is (1(h)), and NG-RC is (1(i)). Only one dimension is displayed (dimension 117).

### v3

@@ V.4 Results
Denoting the true value $\mathbf{y}$, the predicted value $\widehat{\mathbf{y}}$, and the testing set size $h$, the following pointwise error metrics were chosen: the normalized mean squared error (NMSE) given by
$\textup{NMSE}(\mathbf{y},\widehat{\mathbf{y}})=\frac{1}{d}\sum_{u=1}^{d}\frac{\sum_{i=1}^{h}(y_{u,i}-\widehat{y}_{u,i})^{2}}{\sum_{i=1}^{h}(y_{u,i}-\bar{y}_{u})^{2}},$
where $\bar{y}_{u}$ denotes the average of the $u$-th dimension of the vector $\mathbf{y}$ over time steps $i=1,\ldots,h$, the mean absolute error (MAE) given by
$\textup{MAE}(\mathbf{y},\widehat{\mathbf{y}})=\frac{1}{h}\sum_{i=1}^{h}\|\mathbf{y}_{i}-\widehat{\mathbf{y}_{i}}\|_{1}$
where $\|\cdot\|_{1}$ is the 1-norm, the median absolute error (MdAE) defined as
$\textup{MdAE}(\mathbf{y},\widehat{\mathbf{y}})=\frac{1}{d}\sum_{u=1}^{d}\textup{median}\left(\{|y_{u,i}-\widehat{y}_{u,i}|\}_{i=1}^{h}\right),$
and the mean absolute percentage error (MAPE)
@@
for some very small $\varepsilon>0$.
Whether the estimator replicated the climate of the true time series was measured by considering the difference in the true and estimated power spectral density (PSD) and the difference in distributions using the Wasserstein-1 distance.
The PSD is the Fourier transform of the autocovariance function and represents the time series in its frequency domain [50]. The PSD of each time series was estimated by periodograms computed using Welch’s method, provided by scipy.signal.welch, with the Hann window. The number of points per segment was chosen by visual inspection for a balance between frequency resolution and error variance. When the PSD tapers off to zero after some frequency $F$, the remaining frequencies are not considered in the final difference. Finally, the PSD error (PSDE) is computed by taking
@@
The Wasserstein-1 distance or earth mover distance, arising out of optimal transport [51], is a measure of the distance between probability measures and is given by
@@
where $\Gamma(\mu,\widehat{\mu})$ is the set of probability distributions whose marginals are $\mu$ and $\widehat{\mu}$ on the first and second factors respectively. $\mu$ and $\widehat{\mu}$ are the joint distributions of $\mathbf{y}$ and $\widehat{\mathbf{y}}$ respectively.
The Wasserstein-1 distance was computed to compare the distributions of the true and estimated systems in our paper. For one-dimensional systems, the Wasserstein-1 distance can be computed using scipy.stats.wasserstein_distance which uses the equivalent Cramer-1 distance, that is,
@@
where $\mu,\widehat{\mu}$ are the probability distributions of $y$ and $\widehat{y}$ respectively while CDF, $\widehat{\textup{CDF}}$ denote the cumulative distributive functions. For time series in $d$-dimensions, using scipy.stats.wasserstein_distance_nd, corresponds to solving the linear programming problem in (25). It is noted that computing the Wasserstein distance for the multidimensional case is significantly more computationally expensive than in the one-dimensional case. In the case of the Lorenz dynamical system, sampling had to be performed to make using scipy.stats.wasserstein_distance_nd tractable.
@@
The performance of the estimators was measured in the following manner for the dynamical systems. First, the valid prediction time was computed. Then, the ceiling of the best-performing valid prediction time is taken. The point-to-point error metrics NMSE, MAE, MdAE, and MAPE are measured only up to $\lceil T_{\textup{valid}}\rceil$. Beyond the characteristic predictable timescale given by the Lyapunov time, it is not meaningful to measure the step-to-step error as the trajectories diverge exponentially according to the Lyapunov exponent. To determine if the climate is well replicated over the full testing dataset, the climate metrics PSDE and $W_{1}$ are computed for the full testing dataset, with the Lorenz needing to be sampled to compute $W_{1}$. The errors are reported in Table 3.
Table 3: Error values for each estimator and dataset. A 20% deviation is allowed for $T_{\textup{valid}}$. NMSE, MAE, MdAE, and MAPE are computed up to $\lceil T_{\textup{valid}}\rceil$ for Lorenz and Mackey-Glass. PSDE and $W_{1}$ are computed for the full dataset (Lorenz was sampled). PSDE for BEKK is scaled by $10^{-4}$.
System Estimator $T_{\textup{valid}}$ NMSE MAE MdAE MAPE PSDE $W_{1}$
Lorenz NG-RC 7.178 0.379 1.827 0.0897 1.598 7.606 2.114
Polynomial 9.126 0.188 1.155 0.0377 1.230 7.169 1.683
Volterra 10.566 0.0934 0.428 0.00385 0.222 8.325 1.982
Mackey-Glass NG-RC 0.3 0.980 0.188 0.174 0.235 28.815 0.155
Polynomial 7.035 0.0699 0.0274 0.00751 0.0329 6.831 0.00147
Volterra 8.305 0.0333 0.0162 0.00202 0.0190 5.059 0.00138
BEKK NG-RC 0.875 0.0204 0.0166 0.644 1.565 0.332
Polynomial 0.877 0.0204 0.0166 0.642 1.140 0.337
Volterra 0.619 0.0170 0.0139 0.634 0.963 0.319
@@
ct-.15inVolterra Refer to caption
@@
ct-.15inPolynomial Refer to caption
@@
ct-.15inNG-RC Refer to caption
@@
Figure 1: The PSD for: (1(a)) the Lorenz system for the Volterra kernel, (1(b)) the polynomial kernel, and (1(c)) NG-RC. Dimension $1$ corresponds to $x$, dimension $2$ to $y$, and dimension $3$ to $z$. PSD for Mackey-Glass system for Volterra kernel is (1(d)), polynomial kernel is (1(e)), and NG-RC is (1(f)). Dimension 1 refers to the $z$ value. PSD for BEKK for Volterra kernel is (1(g)), polynomial kernel is (1(h)), and NG-RC is (1(i)). Only the two dimensions with the most visually prominent PSD are displayed (dimensions 85 and 118).
@@
ct-0.15inVolterra Refer to caption
@@
ct-0.15inPolynomial Refer to caption
@@
ct-0.15inNG-RC Refer to caption
@@
Figure 2: The distributions for Lorenz system for the Volterra kernel is in (1(a)), the polynomial kernel is in (1(b)), and NG-RC is in (1(c)). The distribution for the Mackey-Glass system for Volterra kernel is (1(d)), the polynomial kernel is (1(e)), and NG-RC is (1(f)). The distributions for BEKK for Volterra kernel is (1(g)), the polynomial kernel is (1(h)), and NG-RC is (1(i)). Only one dimension is displayed (dimension 118).
@@
Overall, we find that the numerical simulations illustrate the points made in the theory discussed above. First, since the optimization problem in the NG-RC methodology is equivalent to that solved in polynomial kernel regression (as per Proposition 3), in more complex systems, it is better to kernelize to access computationally a richer feature space. This is illustrated especially by the superior performance of the polynomial kernel regression against the NG-RC in the Mackey-Glass system. Second, using the universal Volterra kernel, one can consider infinite lag and monomial powers, which is especially advantageous when learning significantly more complex systems, such as the BEKK input/output systems. In such cases, limited to finite lags and monomial orders, which need to be chosen (the Volterra kernel is agnostic to them), the polynomial kernel regression and NG-RC methodologies fail.